OpenStudy (anonymous):

Show that.. Integral of 0 to 6 (1/(x+2))dx = 2ln2

OpenStudy (anonymous):

\[\int\limits_{0}^{6}(1/(x+2))dx = 2\ln2\]

OpenStudy (anonymous):

\[\int\limits_{}^{} 1/(x+2) dx = \ln(x+2)\] (I am excluding the constant C since we are dealing with a definite integral) So now we take the limit from 0 to 6 using the Fundamental Theorem of Calculus and get \[\ln 8 -\ln 2\]. This is the same as \[\ln 2^3 - \ln 2\] Do you remember your properties for log functions?

OpenStudy (anonymous):

Not too well. :/ I need to review them again.

OpenStudy (anonymous):

Ok well \[\ln a^b = b*\ln a\] Thus \[\ln 2^3 = 3\ln 2\] so 3ln2 - ln 2 = 2ln2

OpenStudy (anonymous):

ah ok. :D Thanks. Could you please explain to me how you got ln8 - ln2 at first though?

OpenStudy (anonymous):

Suppose \[\int\limits_{}^{} fdx = F\] The Fundamental Theorem of Calculus says \[\int\limits_{a}^{b} fdx = F(b) - F(a)\] In this case \[1/(x+2) = f\] and \[\ln(x+2) = F\] From there it's just plug n chug

OpenStudy (anonymous):

Oh, and a = 0, b = 6 of course

OpenStudy (anonymous):

Thank youuu